# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ 4.2. $L_{p}$ estimates for inhomogeneous systems
Let $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},r>0$, $\lambda\geq 0$, and $\nu\in[32,\infty)$. Suppose that $G=|F|+|f|\in L_{2}(Q_{\nu r}^{+}(z_{0}),\mu)$ and $u\in\mathcal{H}^{1}_{2}(Q_{\nu r}^{+}(z_{0}),\mu)$ is a weak solution to (4.1)-(4.2) in $Q_{\nu r}^{+}(z_{0})$. Then, we can write $u(t,x)=v(t,x)+w(t,x)$ in $Q_{\nu r}^{+}(z_{0})$, where $v$ and $w$ are in $\mathcal{H}^{1}_{2}(Q_{\nu r}^{+}(z_{0}),\mu)$ and satisfy
@@
$\fint_{Q_{r}^{+}(z_{0})}|D_{x^{\prime}}w-(D_{x^{\prime}}w)_{Q_{r}^{+}(z_{0})}|d\mu+\fint_{Q_{r}^{+}(z_{0})}|\mathcal{W}-(\mathcal{W})_{Q_{r}^{+}(z_{0})}|d\mu$ (4.14)
$\,\,+\sqrt{\lambda}\fint_{Q_{r}^{+}(z_{0})}|w-(w)_{Q_{r}^{+}(z_{0})}|d\mu\leq N\nu^{-1}\left(\fint_{Q_{\nu r}^{+}(z_{0})}|U|^{2}d\mu\right)^{1/2}+N\nu^{-1}\left(\fint_{Q_{\nu r}^{+}(z_{0})}|G|^{2}d\mu\right)^{1/2},$
where $N=N(d,n,\alpha,\kappa)>0$ and
@@
We may assume that $\lambda>0$. The case when $\lambda=0$ follows by adding $\eta u$ with $\eta>0$ to both sides of the equation and then taking the limit as $\eta\to 0$. Let $v\in\mathcal{H}^{1}_{2}(\Omega_{t_{0}},\mu)$ be the weak solution to
@@
Let w=uvw=u-v. Then $w\in\mathcal{H}^{1}_{2}(Q_{\nu r}^{+}(z_{0}),\mu)$ is a weak solution to the homogeneous system
@@
Let $\lambda>0$ and $F,f\in L_{\infty}(\mathbb{R}^{d+1}_{+})$ have bounded supports. Let $u\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$ be the weak solution to (4.1)-(4.2) in $\mathbb{R}^{d+1}_{+}$. Then $u\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$ for any $p\in(1,2)$.
@@
Let $(R_{k})_{k=0}^{\infty}$ be an increasing sequence of positive numbers with $R_{k}\to\infty$ as $k\to\infty$, which is to be specified later. Denote $\hat{Q}_{k}=(-R_{k}^{2},R_{k}^{2})\times D_{R_{k}}^{+}$. Fix $\eta_{k}\in C^{\infty}(\mathbb{R}^{d+1})$ such that $\eta_{k}=1$ on $\mathbb{R}^{d+1}_{+}\setminus\hat{Q}_{k+1}$, $\eta_{k}=0$ on $\hat{Q}_{k}$, $|D\eta_{k}|\leq\frac{N_{d}}{R_{k+1}-R_{k}}$, and $|\partial_{t}\eta_{k}|\leq\frac{N_{d}}{(R_{k+1}-R_{k})^{2}}$. Take $R_{0}$ large enough so that $\operatorname{supp}F\cup\operatorname{supp}f\subset\overline{\hat{Q}_{0}}$.
@@
Let $\lambda>0$ and $F,f\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+})$. Let $u\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$ be the weak solution to (4.1)-(4.2) in $\mathbb{R}^{d+1}_{+}$. Then $u,Du\in L_{p}(\mathbb{R}^{d+1}_{+},\mu)$ for any $p\in(2,\infty]$.
@@
By interpolation, it is sufficient to prove the claim for $p=\infty$. With $F,f\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+})$, we can essentially repeat the proofs of Lemmas 4.4 and 4.5, and therefore,
@@
The Hardy-Littlewood and Fefferman-Stein maximal functions of a function $f$ on $\Omega_{T}$ with respect to measure $\mu$ are respectively defined by
@@
Let $\alpha\in(-1,\infty),p,q\in(1,\infty)$, and $w=w_{0}(t)w_{1}(x)$, where $w_{0}\in A_{q}(\mathbb{R})$ and $w_{1}\in A_{p}(\mathbb{R}_{+},\mu)$. Then for any $f\in L_{q,p}(\Omega_{T},wd\mu)$, it holds that
@@
where we used $\mu(Q_{\nu r}^{+}(z_{0}))\leq N_{d,\alpha}\nu^{d+2+\alpha_{+}}\mu(Q_{r}^{+}(z_{0}))$. It now follows that on $\Omega_{T}$, we have
$(D_{x^{\prime}}u)^{\sharp}+(\mathcal{U})^{\sharp}+(\sqrt{\lambda}u)^{\sharp}\leq N\nu^{-1}\mathcal{M}^{1/2}(|U|^{2})$
$\quad+N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{2}})(\mathcal{M}^{1/2}(|F|^{2})+\mathcal{M}^{1/2}(|f|^{2})).$
@@
$+N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{2}})(\|F\|_{L_{p}(\Omega_{T},\mu)}+\|f\|_{L_{p}(\Omega_{T},\mu)}).$
@@
To prove the existence of solutions with $p\in(2,\infty)$, assume without loss that $T=\infty$, that is $\Omega_{T}=\mathbb{R}^{d+1}_{+}$, and take sequences $F_{k},f_{k}\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+})$ such that $F_{k}\to F$ and $f_{k}\to f$ in $L_{p}(\mathbb{R}^{d+1}_{+},\mu)$. Then by the $L_{2}$ solvability, there exist solutions $u_{k}\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$ to (4.1)-(4.2) with $F_{k}$ and $f_{k}$ in place of $F$ and $f$. By Lemma 4.12, $u_{k}\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$. By the a-priori estimate (4.3), the sequence $u_{k}$ is Cauchy in $\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$, and therefore converges to the solution $u\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$ with $F,f$.
For $p\in(1,2)$, we use a duality argument. Let $G\in L_{p^{\prime}}(\Omega_{T},\mu)^{d}$ and $g\in L_{p^{\prime}}(\Omega_{T},\mu)$ be arbitrary. If $T<\infty$, we also extend $G$ and $g$ to be zero for $t\geq T$. As $p^{\prime}\in(2,\infty)$, by reversing the time variable, there exists a unique weak solution $v\in\mathcal{H}^{1}_{p^{\prime}}(\mathbb{R}^{d+1}_{+},\mu)$ to the adjoint equation
@@
Here one can use the Steklov averages to justify taking $u$ and $v$ as test functions in the weak formulation of solutions. We have
@@
To prove the existence of solutions with $p\in(1,2)$, again assume without loss that $T=\infty$, and take sequences $F_{k},f_{k}\in L_{\infty}(\mathbb{R}^{d+1}_{+})$ with compact supports such that $F_{k}\to F$ and $f_{k}\to f$ in $L_{p}(\mathbb{R}^{d+1}_{+},\mu)$. Then by the $L_{2}$ solvability, there exist solutions $u_{k}\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu)$ to (4.1)-(4.2) with $F_{k}$ and $f_{k}$ in place of $F$ and $f$. By Lemma 4.11, $u_{k}\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$. By the a-priori estimate (4.3), the sequence $u_{k}$ is Cauchy in $\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$, and therefore converges to the solution $u\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$ to (4.1)-(4.2) with $F,f$. ∎
@@
Let $p\in(1,\infty),z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},r>0$, and $\nu\in[32,\infty)$. Suppose that $G=|F|+|f|\in L_{p}(Q_{\nu r}^{+}(z_{0}),\mu)$ and $u\in\mathcal{H}^{1}_{p}(Q_{\nu r}^{+}(z_{0}),\mu)$ is a weak solution to (4.1)-(4.2) in $Q_{\nu r}^{+}(z_{0})$. Then, we can write $u(t,x)=v(t,x)+w(t,x)$ in $Q_{\nu r}^{+}(z_{0})$, where $v$ and $w$ are in $\mathcal{H}^{1}_{p}(Q_{\nu r}^{+}(z_{0}),\mu)$ and satisfy
@@
where $N=N(d,n,\alpha,p,\kappa)>0$ and $V,U,\mathcal{W}$ are as in Proposition 4.10.
@@ 5. Equations with partially BMO coefficients in weighted spaces
Let $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}}$, $\gamma_{0}\in(0,1],R_{0}\in(0,\infty),\nu\in[32,\infty)$, $r\in(0,\frac{R_{0}}{\nu}]$, and $q_{0},\sigma\in(1,\infty)$. Let $F,f\in L_{q_{0}}(Q_{\nu r}^{+}(z_{0}),\mu)$ and $u\in\mathcal{H}^{1}_{q_{0}\sigma}(Q_{\nu r}^{+}(z_{0}),\mu)$ be a solution to (5.1) in $Q_{\nu r}^{+}(z_{0})$. Then under Assumption 2.1 ($\gamma_{0},R_{0}$) there exists $U^{Q}$ in $Q_{\nu r}^{+}(z_{0})$ such that
$N^{-1}_{d,n,\kappa}U\leq|U^{Q}|\leq N_{d,n,\kappa}U\quad\text{on $Q_{\nu r}^{+}(z_{0})$},$
@@
$\begin{gathered}\fint_{Q_{r}^{+}(z_{0})}|U^{Q}-(U^{Q})_{Q_{r}^{+}(z_{0})}|d\mu\leq N\nu^{\frac{d+2+\alpha_{+}}{q_{0}}}\left(\fint_{Q_{\nu r}^{+}(z_{0})}(|F|^{q_{0}}+|f|^{q_{0}})d\mu\right)^{1/{q_{0}}}\\ +N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}})\left(\fint_{Q_{\nu r}^{+}(z_{0})}|U|^{{q_{0}}\sigma}d\mu\right)^{\frac{1}{{q_{0}}\sigma}},\end{gathered}$ (5.2)
where $N=N(d,n,\alpha,\kappa,q_{0},\sigma)>0$.
@@
$\overline{A}^{ij}(x_{d})=[A^{ij}]_{\nu r,z_{0}^{\prime}}(x_{d}),\quad\overline{C}_{0}(x_{d})=[C_{0}]_{\nu r,z_{0}^{\prime}}(x_{d}).$
Then $u$ satisfies in $Q_{\nu r}^{+}(z_{0})$
@@
$\leq N\Bigl(\fint_{Q_{\nu r}^{+}(z_{0})}(|F|^{q_{0}}+|f|^{q_{0}}+|A^{ij}(z)-\overline{A}^{ij}(x_{d})|^{q_{0}}|Du|^{q_{0}}+|C_{0}(z)-\overline{C}_{0}(x_{d})|^{q_{0}}|\sqrt{\lambda}u|^{q_{0}})d\mu\Bigr)^{1/{q_{0}}}$
@@
$\fint_{Q_{r}^{+}(z_{0})}|D_{x^{\prime}}w-(D_{x^{\prime}}w)_{Q_{r}^{+}(z_{0})}|d\mu+\fint_{Q_{r}^{+}(z_{0})}|\mathcal{W}-(\mathcal{W})_{Q_{r}^{+}(z_{0})}|d\mu+\sqrt{\lambda}\fint_{Q_{r}^{+}(z_{0})}|w-(w)_{Q_{r}^{+}(z_{0})}|d\mu$
@@
Let $U^{Q}=(D_{x^{\prime}}u,\lambda^{1/2}u,\overline{A}^{dj}D_{j}u)$. Combining the estimates above, we get (5.2). ∎
@@
Let $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}}$, $R_{0}\in(0,\infty),\nu\in[32,\infty),r>\frac{R_{0}}{\nu}$, and $q_{0},\sigma\in(1,\infty)$. Suppose $u\in\mathcal{H}^{1}_{q_{0}\sigma}(Q_{\nu r}^{+}(z_{0}),\mu)$ has $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$ for some $s\in\mathbb{R}$ and $\delta\in(0,1)$. Then
@@
$\leq\left(\frac{\delta R_{0}}{r}\right)^{2-\frac{2}{q_{0}\sigma}}\left(\fint_{Q_{r}^{+}(z_{0})}|U|^{{q_{0}}\sigma}d\mu\right)^{\frac{1}{{q_{0}}\sigma}}\leq\delta^{2-\frac{2}{q_{0}\sigma}}\nu^{2-\frac{2}{q_{0}\sigma}}\left(\fint_{Q_{r}^{+}(z_{0})}|U|^{{q_{0}}\sigma}d\mu\right)^{\frac{1}{{q_{0}}\sigma}}.$
@@
We will use the following generalized version of the weighted Fefferman-Stein inequality in the spirit of [37, Theorem 2.7], which is a special case of [13, Corollary 2.8].
@@
Let $\alpha\in(-1,\infty),p,q\in(1,\infty),w=w_{0}(t)w_{1}(x)$, where $w_{0}\in A_{q}(\mathbb{R})$ and $w_{1}\in A_{p}(\mathbb{R}^{d}_{+},\mu)$, and $f,g\in L_{q,p}(\Omega_{T},wd\mu)$. Suppose that for each $m\in\mathbb{Z}$ and $Q^{m}\in\mathbb{C}_{m}^{+}$ there exists a measurable function $f^{C}$ on $Q^{m}$ such that $\frac{1}{K_{3}}|f|\leq f^{C}\leq K_{3}|f|$ for some constant $K_{3}\geq 1$ and
$\fint_{Q^{m}}|f^{C}(z)-(f^{C})_{Q^{m}}|d\mu(z)\leq g(y)\quad\forall y\in Q^{m}.$

### v2

@@ 4.2. $L_{p}$ estimates for inhomogeneous systems
Let $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},r>0$, $\lambda\geq 0$, and $\nu\in[32,\infty)$. Suppose that $G=|F|+|f|\in L_{2}(Q_{\nu r}^{+}(z_{0}),\mu)$ and $u\in\mathcal{H}^{1}_{2}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ is a weak solution to (4.1)-(4.2) in $Q_{\nu r}^{+}(z_{0})$. Then, we can write $u(t,x)=v(t,x)+w(t,x)$ in $Q_{\nu r}^{+}(z_{0})$, where $v$ and $w$ are in $\mathcal{H}^{1}_{2}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ and satisfy
@@
$\fint_{Q_{r}^{+}(z_{0})}|D_{x^{\prime}}w-(D_{x^{\prime}}w)_{Q_{r}^{+}(z_{0}),\mu}|d\mu+\fint_{Q_{r}^{+}(z_{0})}|\mathcal{W}-(\mathcal{W})_{Q_{r}^{+}(z_{0}),\mu}|d\mu$ (4.16)
$\,\,+\sqrt{\lambda}\fint_{Q_{r}^{+}(z_{0})}|w-(w)_{Q_{r}^{+}(z_{0}),\mu}|d\mu\leq N\nu^{-1}\left(\fint_{Q_{\nu r}^{+}(z_{0})}|U|^{2}d\mu\right)^{1/2}+N\nu^{-1}\left(\fint_{Q_{\nu r}^{+}(z_{0})}|G|^{2}d\mu\right)^{1/2},$
where $N=N(d,\alpha,\kappa)>0$ and
@@
We may assume that $\lambda>0$. The case when $\lambda=0$ follows by adding $\eta x_{d}^{\alpha}u$ with $\eta>0$ to both sides of the equation and then taking the limit as $\eta\to 0$. Let $v\in\mathcal{H}^{1}_{2}(\Omega_{t_{0}},\mu;\mathbb{R}^{n})$ be the weak solution to
@@
Let w=uvw=u-v. Then $w\in\mathcal{H}^{1}_{2}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ is a weak solution to the homogeneous system
@@
Let $\lambda>0$ and $F^{1},\dots,F^{d},f\in L_{\infty}(\mathbb{R}^{d+1}_{+};\mathbb{R}^{n})$ have bounded supports. Let $u\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ be the weak solution to (4.1)-(4.2) in $\mathbb{R}^{d+1}_{+}$. Then $u\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$ for any $p\in(1,2)$.
@@
Let $(R_{k})_{k=0}^{\infty}$ be an increasing sequence of positive numbers with $R_{k}\to\infty$ as $k\to\infty$, which is to be specified later. Denote $\hat{Q}_{k}=(-R_{k}^{2},R_{k}^{2})\times D_{R_{k}}^{+}$. Fix $\eta_{k}\in C^{\infty}(\mathbb{R}^{d+1})$ such that $\eta_{k}=1$ on $\mathbb{R}^{d+1}_{+}\setminus\hat{Q}_{k+1}$, $\eta_{k}=0$ on $\hat{Q}_{k}$, $|D\eta_{k}|\leq\frac{N}{R_{k+1}-R_{k}}$, and $|\partial_{t}\eta_{k}|\leq\frac{N}{(R_{k+1}-R_{k})^{2}}$. Take $R_{0}$ large enough so that $\operatorname{supp}F\cup\operatorname{supp}f\subset\overline{\hat{Q}_{0}}$.
@@
Let $\lambda>0$ and $F^{1},\dots,F^{d},f\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+};\mathbb{R}^{n})$. Let $u\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ be the weak solution to (4.1)-(4.2) in $\mathbb{R}^{d+1}_{+}$. Then $u,Du\in L_{p}(\mathbb{R}^{d+1}_{+},\mu)$ for any $p\in(2,\infty]$.
@@
By interpolation, it is sufficient to prove the claim for $p=\infty$. With $F^{i},f\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+};\mathbb{R}^{n})$, we can essentially repeat the proofs of Lemmas 4.4 and 4.5, and therefore,
@@
The Hardy-Littlewood and Fefferman-Stein maximal functions of an, in general, a vector-valued function $f$ on $\Omega_{T}$ with respect to the measure $\mu$ are respectively defined by
@@
Let $\alpha\in(-1,\infty),p,q\in(1,\infty),K\in[1,\infty)$, and $w=w_{0}(t)w_{1}(x)$, where $w_{0}\in A_{q}(\mathbb{R}),w_{1}\in A_{p}(\mathbb{R}^{d}_{+},\mu)$ with $w_{0}]_{A_{q}(\mathbb{R})}\leq K,[w_{1}]_{A_{p}(\mathbb{R}^{d}_{+},\mu)}\leq K$. Then, for any $f\in L_{q,p}(\Omega_{T},wd\mu)$ holds
@@
where we used $\mu(Q_{\nu r}^{+}(z_{0}))\leq N_{\alpha}\nu^{d+2+\alpha_{+}}\mu(Q_{r}^{+}(z_{0}))$. It now follows that on $\Omega_{T}$, we have
$(D_{x^{\prime}}u)^{\sharp}+(\mathcal{U})^{\sharp}+(\sqrt{\lambda}u)^{\sharp}\leq N\nu^{-1}\mathcal{M}^{1/2}(|U|^{2})+N\nu^{\frac{d+2+\alpha_{+}}{2}}(\mathcal{M}^{1/2}(|F|^{2})+\mathcal{M}^{1/2}(|f|^{2})).$
@@
$+N\nu^{\frac{d+2+\alpha_{+}}{2}}(\|F\|_{L_{p}(\Omega_{T},\mu)}+\|f\|_{L_{p}(\Omega_{T},\mu)}).$
@@
To prove the existence of solutions with $p\in(2,\infty)$, assume without loss that $T=\infty$, that is $\Omega_{T}=\mathbb{R}^{d+1}_{+}$, and take sequences $F_{k},f_{k}\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+})$ such that $F_{k}\to F$ and $f_{k}\to f$ in $L_{p}(\mathbb{R}^{d+1}_{+},\mu)$. Then by the $L_{2}$ solvability, there exist solutions $u_{k}\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ to (4.1)-(4.2) with $F_{k}$ and $f_{k}$ in place of $F$ and $f$. By Lemma 4.12, $u_{k}\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$. By the a-priori estimate (4.3), the sequence $u_{k}$ is Cauchy in $\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$, and therefore converges to the solution $u\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ with $F,f$.
For $p\in(1,2)$, we use a duality argument. Let $G\in L_{p^{\prime}}(\Omega_{T},\mu;\mathbb{R}^{n})^{d}$ and $g\in L_{p^{\prime}}(\Omega_{T},\mu;\mathbb{R}^{n})$ be arbitrary. If $T<\infty$, we also extend $G$ and $g$ to be zero for $t\geq T$. As $p^{\prime}\in(2,\infty)$, by reversing the time variable, there exists a unique weak solution $v\in\mathcal{H}^{1}_{p^{\prime}}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ to the adjoint equation
@@
Here one can use Steklov averages to justify taking $u$ and $v$ as test functions in the weak formulation of solutions. We have
@@
To prove the existence of solutions with $p\in(1,2)$, again assume without loss that $T=\infty$, and take sequences $F_{k},f_{k}\in L_{\infty}(\mathbb{R}^{d+1}_{+})$ with compact supports such that $F_{k}\to F$ and $f_{k}\to f$ in $L_{p}(\mathbb{R}^{d+1}_{+},\mu)$. Then by the $L_{2}$ solvability, there exist solutions $u_{k}\in\mathcal{H}^{1}_{2}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ to (4.1)-(4.2) with $F_{k}$ and $f_{k}$ in place of $F$ and $f$. By Lemma 4.11, $u_{k}\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$. By the a-priori estimate (4.3), the sequence $u_{k}$ is Cauchy in $\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu)$, and therefore converges to the solution $u\in\mathcal{H}^{1}_{p}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ to (4.1)-(4.2) with $F,f$. ∎
@@
Let $p\in(1,\infty),z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},r>0$, and $\nu\in[32,\infty)$. Suppose that $G=|F|+|f|\in L_{p}(Q_{\nu r}^{+}(z_{0}),\mu)$ and $u\in\mathcal{H}^{1}_{p}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ is a weak solution to (4.1)-(4.2) in $Q_{\nu r}^{+}(z_{0})$. Then, we can write $u(t,x)=v(t,x)+w(t,x)$ in $Q_{\nu r}^{+}(z_{0})$, where $v$ and $w$ are in $\mathcal{H}^{1}_{p}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ and satisfy
@@
where $N=N(d,\alpha,p,\kappa)>0$ and $V,U,\mathcal{W}$ are as in Proposition 4.10.
@@ 5. Equations with partially BMO coefficients in weighted spaces
Let $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},\gamma_{0}>0,\nu\geq 32,R_{0}\in(0,\infty]$, $r\in(0,\frac{R_{0}}{\nu}]$ $(r\in(0,\infty)$ if $R_{0}=\infty)$, and $q_{0},\sigma\in(1,\infty)$. Let $F^{1},\dots,F^{d},f\in L_{q_{0}}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ and $u\in\mathcal{H}^{1}_{q_{0}\sigma}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ be a solution to (5.1) in $Q_{\nu r}^{+}(z_{0})$. Then under Assumption 2.1 ($\gamma_{0},R_{0}$) there exists $U^{Q}:Q_{\nu r}^{+}(z_{0})\to\mathbb{R}_{\geq 0}$ such that
$N^{-1}_{\kappa}U\leq U^{Q}\leq N_{\kappa}U\quad\text{on $Q_{\nu r}^{+}(z_{0})$},$
@@
$\begin{gathered}\fint_{Q_{r}^{+}(z_{0})}|U^{Q}-(U^{Q})_{Q_{r}^{+}(z_{0}),\mu}|d\mu\leq N\nu^{\frac{d+2+\alpha_{+}}{q_{0}}}\left(\fint_{Q_{\nu r}^{+}(z_{0})}(|F|^{q_{0}}+|f|^{q_{0}})d\mu\right)^{1/{q_{0}}}\\ +N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}})\left(\fint_{Q_{\nu r}^{+}(z_{0})}|U|^{{q_{0}}\sigma}d\mu\right)^{\frac{1}{{q_{0}}\sigma}},\end{gathered}$ (5.2)
where $N=N(d,\alpha,\kappa,q_{0},\sigma)>0$.
@@
$\overline{A}^{ij}(x_{d})=[A^{ij}]_{\nu r,z_{0}}(x_{d}),\quad\overline{C}_{0}(x_{d})=[C_{0}]_{\nu r,z_{0}}(x_{d})$
be from Assumption 2.1 ($\gamma_{0},R_{0}$). Then $u$ satisfies in $Q_{\nu r}^{+}(z_{0})$
@@
$\leq N\Bigl(\fint_{Q_{\nu r}^{+}(z_{0})}(|F|^{q_{0}}+|f|^{q_{0}}+|A(z)-\overline{A}(x_{d})|^{q_{0}}|Du|^{q_{0}}+|C_{0}(z)-\overline{C}_{0}(x_{d})|^{q_{0}}|\sqrt{\lambda}u|^{q_{0}})d\mu\Bigr)^{1/{q_{0}}}$
@@
$\fint_{Q_{r}^{+}(z_{0})}|D_{x^{\prime}}w-(D_{x^{\prime}}w)_{Q_{r}^{+}(z_{0}),\mu}|d\mu+\fint_{Q_{r}^{+}(z_{0})}|\mathcal{W}-(\mathcal{W})_{Q_{r}^{+}(z_{0}),\mu}|d\mu+\sqrt{\lambda}\fint_{Q_{r}^{+}(z_{0})}|w-(w)_{Q_{r}^{+}(z_{0}),\mu}|d\mu$
@@
Let $U^{Q}=|D_{x^{\prime}}u|+\lambda^{1/2}|u|+|\overline{A}^{dj}D_{j}u|$. Combining the estimates above, we get (5.2). ∎
@@
Let $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}}$, $R_{0}\in(0,\infty),\nu\in[32,\infty),r>\frac{R_{0}}{\nu}$, and $q_{0},\sigma\in(1,\infty)$. Suppose $u\in\mathcal{H}^{1}_{q_{0}\sigma}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ has $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$ for some $s\in\mathbb{R}$ and $\delta>0$. Then
@@
$\leq\left(\frac{\delta R_{0}}{r}\right)^{2-\frac{2}{q_{0}\sigma}}\left(\fint_{Q_{r}^{+}(z_{0})}|U|^{{q_{0}}\sigma}d\mu\right)^{\frac{1}{{q_{0}}\sigma}}\leq(\delta\nu)^{2-\frac{2}{q_{0}\sigma}}\left(\fint_{Q_{r}^{+}(z_{0})}|U|^{{q_{0}}\sigma}d\mu\right)^{\frac{1}{{q_{0}}\sigma}}.$
@@
We will use the following generalized version of the weighted Fefferman-Stein inequality in the spirit of [41, Theorem 2.7], which is a special case of [14, Corollary 2.8].
@@
Let $\alpha\in(-1,\infty),p,q\in(1,\infty),K,K_{4}\in[1,\infty)$, $w=w_{0}(t)w_{1}(x)$, where $w_{0}\in A_{q}(\mathbb{R}),w_{1}\in A_{p}(\mathbb{R}^{d}_{+},\mu)$ with $w_{0}]_{A_{q}(\mathbb{R})}\leq K,[w_{1}]_{A_{p}(\mathbb{R}^{d}_{+},\mu)}\leq K$, and $f,g\in L_{q,p}(\Omega_{T},wd\mu)$. Suppose that for each $m\in\mathbb{Z}$ and $Q^{m}\in\mathbb{C}_{m}^{+}$ there exists a measurable function $f^{C}$ on $Q^{m}$ such that $\frac{1}{K_{4}}|f|\leq f^{C}\leq K_{4}|f|$ a.e. on $Q^{m}$, and
$\fint_{Q^{m}}|f^{C}(z)-(f^{C})_{Q^{m},\mu}|d\mu(z)\leq g(y)\quad\forall y\in Q^{m}.$